OpenStudy (anonymous):

find x in the equation and assume that any segment that appears to be tangent is a tangent theta=x small=45

OpenStudy (anonymous):



Directrix (directrix):

small=45 --> I don't know what you mean by this. ========================== The formula:the angle formed by the intersection of 2 tangents outside the circle equals half the difference of the intercepted arcs. 34 = (1/2) [ (360 -x) - x] 68 = 360 - 2x x = 146

OpenStudy (anonymous):

Thank you